# A systematic comparison of measures for k-anonymity in networks

### v1

@@ 3.2 Intermediate representation
The second category we discuss is an intermediate representation of the network. Two forms of intermediate representations have been introduced for privacy aware sharing of networks, being clustering and injecting uncertainty [32, 27, 33, 55].
For the clustering approaches the nodes of the network are first merged into super nodes [32, 27, 33, 34, 35, 36, 37] according to some predefined mechanism. These supernodes additionally contain labels denoting the number of nodes and edges within each cluster. If there is at least one edge between nodes in two clusters, this is modelled using a superedge denoting the number of connections between nodes in the different clusters. The resulting graph can then be used to generate a network to analyze. By ensuring each cluster has a size of at least $k$, anonymity of the entities can be ensured. Due to the clustering, this approach likely preserves global properties, yet local properties may be destroyed as the exact neighborhood structure surrounding a node can not be inferred from the supernodes and superedges. Different works focus on improving the clusters made [35, 36], and some approaches add additional constraints to account for node labels [37].
Another approach is to inject uncertainty into edges. The work presented in [55] introduces an approach that adds a probability to each possible edge in the graph to hide certain node properties. This resulting graph with edge probabilities can be shared and used to sample new synthetic networks. This approach would hence ensure edge anonymity. However, which network properties would be preserved and destroyed depends heavily on how the probability is assigned to the edges.
3.3 Synthetic data
The third category of approaches generates a synthetic network based on graph properties of the original network. Various generic methods exist to generate network models that aim to capture real-world network properties. Examples are the Barabási–Albert [59] or Watts-Strogatz model [60]. An example of such a model specifically for achieving privacy, is the dK-graph model [58]. However, it is found that this in itself may not ensure privacy as many nodes can still be identified in the resulting dK-graph [61]. Hence, subsequentially introduced work first generates a differentially private joint degree distribution, as explained in Section 3.1, which can then be used to generate a dK-graph [25, 54].
@@ 3.4 Perturbed network
The last category is perturbation. For this type of approach, the goal is to share an adapted version of the original network which does not reveal a node’s identity based on structure. Overall there are two types of approaches: randomization [29, 30, 31] to ensure edge privacy, and $k$-anonymity [26, 27, 11, 28] to ensure node privacy.
Various operations on edges, such as edge addition, deletion or rewiring, can be used to perturb the network. After the graph is perturbed, an attacker may not be certain about the existence of a connection between two entities, hence ensuring edge anonymity. The work in [29] introduces such an approach, additionally preserving utility by retaining the spectrum, i.e., the set of eigenvalues of the adjacency matrix representing the network. In [30, 31], edges are added based on random walks with the aim of preserving more data utility.
One of the most common methods for ensuring node privacy, which is also commonly used in SDC literature and the focus of the remainder of this paper, is $k$-anonymity [26, 27, 11, 28]. When using this approach, for each node in the graph there should be $k$ equivalent candidates given an amount of structural information hence ensuring node privacy. This amount of information, i.e., the chosen anonymity measure, corresponds to an attacker scenario where we assume that a possible attacker has this amount of knowledge. Based on this measure we can determine if a given node is $k$-anonymous or not. When choosing the measure, the privacy versus utility trade-off is very important, as discussed in Section 2.
Different anonymization algorithms can be used which could aim to preserve certain topological properties. Inevitably, some properties need to be destroyed in order to make all nodes $k$-anonymous. An example is that perturbation will always change the degree distribution, even if the chosen anonymity measure is only the degree of nodes. In the remainder of this paper we focus on $k$-anonymity by surveying and comparing different measures.
4 An overview of k-anonymity measures
In this section, we first define the required network concepts and notions related to $k$-anonymity. Second, we provide an overview of measures for $k$-anonymity introduced in literature thus far.
@@ 4.1 Networks
We define a network or graph $G=(V,E)$ as a set of nodes $V$, and a set of edges $\{v,w\}\in E$, with $v,w\in V$. The degree of a node equals the number of connections it has: $degree(v)=|\{w:\{v,w\}\in E\}|$. We can create a distribution of node degrees in the network. In real-world networks, this often resembles a powerlaw distribution [62] where $occurrence(degree)=c*degree^{-\alpha}$. Here $c$ is a constant and $\alpha$ the powerlaw exponent. The value for $\alpha$ indicates the steepness of the slope of the degree distribution. Hence, a low $\alpha$ indicates a less steep slope and a more diverse degree distribution, while a high value implies a more steep slope and a less diverse degree distribution.
@@
In real-world networks it often occurs that two neighboring nodes form a triangle with a third node. We measure the tendency to form triangles using the clustering coefficient. For a node, this equals the number of triangles it is part of, divided by the maximum number of triangles it could have which equals $\frac{1}{2}degree(v)(degree(v)-1)$. For a graph, we summarize the clustering coefficient as the average clustering coefficient over all nodes.
Nodes may connect to nodes with a similar degree or a different degree. We capture this tendency by assortativity. A negative value implies nodes tend to connect to dissimilar nodes while a positive value indicates they tend to connect to similar nodes.
@@
We define the distance between two nodes, $distance(v,w)$, as the minimum number of edges that needs to be traversed to reach one node from the other. It follows that $distance(v,w)=distance(w,v)$ and $distance(v,v)=0$. When there is no path between two given nodes, i.e., a sequence of edges connecting $v$ and $w$, then $distance(v,w)=\infty$. For simplicity, we focus on undirected graphs. This occurs when the nodes are in different components, i.e., maximal subsets of nodes in which there is a path between all pairs of nodes. The diameter of the graph, $D(G)$, is the length of the longest shortest path between two nodes that is not equal to $\infty$. The average distance equals the average of the length of all shortest paths between all node pairs in the network that are in the same component.
We define the $d$-neighborhood of a node $N_{d}(v)=(V_{N_{d}(v)},E_{N_{d}(v)})$ as the set of nodes that are at most distance $d$ from target node $v$, and set of all edges between these nodes. Two neighborhoods in a graph are structurally indistinguishable if they are isomorphic, as defined in Definition 1.
Definition 1 (Graph isomorphism)
Given two graphs $G=(V,E)$ and $G^{\prime}=(V^{\prime},E^{\prime})$, a graph isomorphism is a bijective function $\phi:V\rightarrow V^{\prime}$ such that for each $v,w\in V$ it holds that $\{\phi(v),\phi(w)\}\in E^{\prime}$ precisely when $\{v,w\}\in E$.
@@
We can determine if two graphs are isomorphic by comparing their canonical labeling [63]. This is a label assigned by a function $\mathcal{C}$ such that two graphs have the same label value only if they are isomorphic. A special case of isomorphism is automorphism: an isomorphism from the graph onto itself. If two nodes are mapped onto each other by an automorphism, they are in the same orbit. This implies that the nodes are structurally indistinguishable from each other.
@@
In order for a node to be $k$-anonymous, it should be equivalent to at least $k-1$ other nodes according to a particular definition for equivalence $M$. We refer to $M$ as the anonymity measure. If two nodes $v,w\in V$ are equivalent using a measure $M$, we denote this as $v\cong_{M}w$. These nodes are said to be in the same equivalence class. We let $EC_{M}$ denote the partitioning of nodes $V$ into equivalence classes and $EC_{M}(v)$ the equivalence class of a given node $v$. We say that a node $v$ is $k$-anonymous when $|EC_{M}(v)|=k$. We summarize the anonymity of a graph given a certain measure by its uniqueness, defined in Definition 4.
Definition 2 (Equivalence class)
Given a graph $G=(V,E)$ and anonymity measure $M$, we define an equivalence class $EC_{M}$ as a set of nodes, such that $\forall v,w\in EC_{M}:v\cong_{M}w$.
Definition 3 (k-Anonymity)
Given a graph $G=(V,E)$ and measure $M$ a node is $k$-anonymous for $M$ if it is in an equivalence class of size $k$. The graph $G$ is $k$-anonymous if all nodes in $V$ are at least $k$-anonymous.
Definition 4 (Uniqueness)
Given a graph $G=(V,E)$, measure $M$ and equivalence classes $EC_{M}$, a node is unique if $|EC_{M}(v)|=1$. We define the uniqueness of the graph as the fraction of unique nodes in the graph.
$uniqueness(G)=\frac{|\{v:v\in V,|EC_{M}(v)|=1\}|}{|V|}$
4.3 k-Anonymity measures
In this section, we discuss measures for $k$-anonymity introduced in literature. We categorize these based on what type of information they take into account: degree based, neighborhood based, automorphism based and hybrid measures. While the selection of a measure can be done separately from the anonymization process, the meausure does influence the difficulty of anonymizing a graph. Hence, for each measure we both describe the measure itself and give an estimate of how complex it is to achieve anonymity with as few alterations as possible. A parameterized version for most measures can be created to account for structural information beyond the direct neighborhood of the target node, denoted with parameter $d$. Additionally, we discuss how each of the discussed measures can be extended by a cascading step.
We summarize the measures we focus on throughout the remainder of the manuscript in Table 2. These measures are chosen based on two criteria. First, each of the measures assigns one or more values to each node, which are used to divide the set of all nodes into equivalence classes. Second, the chosen measures represent the set of measures introduced in literature that model feasible attacker scenarios.
@@ 4.3.1 Degree based
We distinguish between three measures within the category of degree based measures, being degree, vertex refinement queries and joint degree.
@@
Most of the works using degree based measures focus on the actual degree of the node [26, 68, 69, 70, 64, 43, 44]. A graph is then $k$-anonymous if each degree occurs at least $k$ times in the network. Compared to the other measures we discuss, this is the simplest measure, as it accounts for the least structural information and is trivially easy to compute for a given node. As a result, anonymization algorithms have been introduced that minimize the number of edge deletions. These approaches can be used on networks with over a million nodes and edges [26, 43, 44]. Other works that introduced anonymization methods for this measure specifically aim to preserve more data utility than the other approaches [68, 70, 64].
@@
A parameterized measure that takes into account the degree of nearby nodes is introduced in [38, 27]. This work introduces the notion of vertex refinement queries (VRQ), denoted $\mathcal{H}_{i}$, for which the distance can be set with parameter $i$. When $\mathcal{H}_{1}$ is used, only the degree of the node is taken into account. For $\mathcal{H}_{i}$, with $i>1$, the degree distribution of the nodes at distance at most $i-1$ is taken into account. The measure is applied recursively, i.e., $\mathcal{H}_{i}$ is used to compute $\mathcal{H}_{i+1}$ for each $i\geq 1$. This measure is more difficult to anonymize for and requires an approximation algorithm minimize the number of edge deletions.
@@
A measure that is more strict than degree, but not as strict as VRQ, has been introduced in [71] and accounts for pairs of degrees: the degree of the target node and its neighbor. While this measure accounts for only slightly more information than node degree itself, anonymization is already much more difficult. To anonymize a network given this measure, the work proposes an integer programming problem and an approximation algorithm to anonymize larger networks. This measure differs from the remaining measures as it does not compute a value for a node to determine the equivalence classes. Hence we do not include the measure in our comparison. However, it does show similarities to the cascading algorithm we discuss in Section 4.3.5 and 6.4.
4.3.2 Neighborhood based
@@
For neighborhood based measures, we distinguish between measures that account for complete structural information and measures that are based on meaningful properties of the neighborhood. Most common are measures that account for the complete structure of the 1-neighborhood [11, 45, 51, 72, 73, 66]. Here nodes are said to be equivalent if their neighborhoods are isomorphic, which implies the neighborhoods are not distinguishable based on structure. Other works propose a parameterized version of this measure assuming structural knowledge of the $d$-neighborhoods [66, 73, 65, 12] resulting in a measure referred to as $d$-$k$-anonymity. Similar to VRQ, this measure is computed recursively. Hence, in order for a node to be $d$-$k$-anonymous, it needs to be $(d-1)$-$k$-anonymous.
@@
The work in [65] uses heuristics to speed up the computation of $d$-$k$-anonymity. These properties are graph invariants: properties that are equal if two graphs, or in this case $d$-neighborhoods, are isomorphic. If the values are not equal, the $d$-neighborhoods can not be isomorphic and as a result the nodes can not be equivalent. However, the converse does not hold: if the graph invariants are equal it does not imply that the neighborhoods are isomorphic. The heuristics used are count, which equals the number of nodes and edges in the $d$-neighborhood and degree, which equals the degree distribution of the $d$-neighborhood. From the degree distribution the number of nodes and edges in the $d$-neighborhood, which equals the value for count, can be derived. As these heuristics in themselves can be used as measures for equivalence and are in terms of what information they measure elegantly positioned between degree and $d$-$k$-anonymity, we choose to include these in our comparison.
4.3.3 Automorphism based
Several works focus on a very strict scenario where an attacker should not be able to distinguish between two nodes, even with complete structural information [28, 39, 74]. There are two variants. In [28, 39], the aim is that for each node there should be at least $k-1$ nodes in the same orbit. This implies that the nodes have the exact same structural properties and hence can not be distinguished even when an attacker has perfect structural information over the entire network. For [74], the network is $k$-anonymous if it can be partitioned into $k$ subgraphs that are isomorphic to each other. To achieve $k$-anonymity for these measures, symmetry should be introduced to the network. In particular, for a network to be $k$-anonymous, all its components should be symmetric in at least $k-1$ points, which is likely not realistic for most real-world networks. This would moreover have major impact on the utility of an anonymized version of the network. As this approach is computationally expensive and approximated by the distance-parameterized structural neighborhood measures mentioned at the end of Section 4.3.2, we do not include automorphism based measures in our comparison.
4.3.4 Hybrid

### v2

@@ 2.3 Intermediate representation
The third category we discuss outputs an intermediate representation of the original network. This intermediate representation can be directly analyzed, or a network can be sampled from the representation. Two types of intermediate representations have been introduced for the purpose of sharing privacy sensitive network data, being clustering [37, 32, 38, 39, 40, 41] and injecting uncertainty [51].
For the clustering approaches the nodes of the network are merged into super nodes [37, 32, 38, 39, 40, 41] according to some predefined mechanism. These supernodes additionally contain labels denoting the number of nodes and edges within each cluster. If there is at least one edge between nodes in two clusters, this is modeled using a superedge denoting the number of connections between nodes in the different clusters. Various works focus on improving the clusters made [40], and some approaches add additional constraints to account for node labels [41].
By ensuring that each cluster has a size of at least a preset number of nodes, anonymity of the entities can be ensured. The clustering process can to some extent preserve global properties, yet local properties may be destroyed as the exact neighborhood structure surrounding a node can not be inferred from the supernodes and superedges.
The second approach in this category is to add uncertainty by assigning a probability to each edge in the network to hide certain node properties [51]. The resulting graph with edge probabilities can be published and used to sample new networks. This approach would hence ensure edge anonymity. Which network properties would be preserved and destroyed depends on how the probability is assigned to the edges.
@@ 2.4 Perturbed network
For the last category of approaches, the goal is to share an adapted version of the original network such that privacy of entities represented is ensured. Overall there are two types of approaches: randomization [34, 35, 36] to ensure edge privacy, and $k$-anonymity [31, 32, 11, 33] to ensure node privacy.
To protect edge privacy, operations on edges including edge addition, deletion or rewiring, can be used to perturb the network. After the graph is perturbed, an attacker can not be certain about the existence of a connection between two entities, hence ensuring edge anonymity. The work in [34] introduces such an approach, additionally preserving utility by retaining the spectrum, i.e., the set of eigenvalues of the adjacency matrix representing the network. In [35, 36], edges are added based on random walks with the aim of preserving more data utility.
One of the most common methods for ensuring node privacy, which is also the focus of this work, is $k$-anonymity [31, 32, 11, 33]. An anonymity measure is used to assess whether a graph satisfies $k$-anonymity, which is true when for each node in the graph there are at least $k-1$ equivalent candidates given an amount of structural information indicated by the anonymity measure. The chosen anonymity measure corresponds to an attacker scenario where we assume that a possible attacker has this amount of knowledge. If the network is not $k$-anonymous an anonymization algorithm can be applied. These algorithms can be designed to preserve utility, i.e., certain structural properties or performance on downstream tasks [42]. Inevitably, some properties need to be destroyed in order to make all nodes $k$-anonymous.
In the remainder of this paper, we focus on $k$-anonymity as this approach allows one to safely publish or share an altered version of the network.
3 Preliminaries
In this section, we define the required concepts on networks and $k$-anonymity.
@@ 3.1 Networks
We define a network or graph $G=(V,E)$ as a set of nodes $V$, and a set of edges $\{v,w\}\in E$, with $v,w\in V$. The degree of a node equals the number of connections it has: $degree(v)=|\{w:\{v,w\}\in E\}|$. We can create a distribution of node degrees in the network. In real-world networks, this often resembles a powerlaw distribution [59] where degrees are distributed as $c*degree^{-\alpha}$. Here $c$ is a constant and $\alpha$ the powerlaw exponent. The value of $\alpha$ indicates the steepness of the slope of the degree distribution. Hence, a lower value of $\alpha$ indicates a heavier tail and greater imbalance between low- and high-degree nodes, whereas a higher $\alpha$ results in a steeper slope, implying a more balanced degree distribution with fewer high-degree nodes.
@@
In real-world networks it often occurs that two neighboring nodes form a triangle with a third node. We measure the tendency to form triangles using the node clustering coefficient. For a node, this equals the number of triangles it is part of, divided by the maximum number of triangles it could be part of, which equals $\frac{1}{2}degree(v)(degree(v)-1)$. For a graph, we summarize the clustering coefficient as the average clustering coefficient over all nodes. Nodes may connect to nodes with a similar degree or a different degree. We capture this tendency by assortativity. A negative value means nodes tend to connect to dissimilar nodes while a positive value indicates they tend to connect to similar nodes.
@@
We define the distance between two nodes, $distance(v,w)$, as the minimum number of edges that needs to be traversed to reach one node from the other. Since the edges are undirected, it follows that $distance(v,w)=distance(w,v)$ and $distance(v,v)=0$. When there is no path between two given nodes, i.e., a sequence of edges connecting $v$ and $w$, then $distance(v,w)=\infty$. This occurs when the nodes are in different components, i.e., maximal subsets of nodes in which there is a path between all pairs of nodes. The diameter of the graph, $D(G)$, is the largest shortest path length between two nodes that is not equal to $\infty$. The average distance in a graph equals the average of the length of all shortest paths between all node pairs in the network that are in the same component.
@@
We define the $d$-neighborhood of a node $N_{d}(v)=(V_{N_{d}(v)},E_{N_{d}(v)})$ as the set of nodes that are at most distance $d$ from the considered node $v$, and the set of all edges between these nodes. When two neighborhoods in a graph are structurally indistinguishable they are isomorphic. More precisely, two graphs $G=(V,E)$ and $G^{\prime}=(V^{\prime},E^{\prime})$ are isomorphic if there exists a bijective function $\phi:V\rightarrow V^{\prime}$ such that for each $v,w\in V$ it holds that $\{\phi(v),\phi(w)\}\in E^{\prime}$ precisely when $\{v,w\}\in E$. We can determine if two graphs are isomorphic by comparing their canonical labeling [60]. This is a label assigned by a function $\mathcal{C}$ such that two graphs have the same label value only if they are isomorphic. A special case of isomorphism is automorphism: an isomorphism from the graph onto itself. If two nodes can be mapped onto each other by an automorphism, they are in the same orbit. This implies that the nodes are structurally indistinguishable from each other.
To assess the utility of the anonymized networks in Section 6.4, we use three metrics that are commonly used in downstream network analysis tasks. Most of the nodes in real-world networks are in the largest connected component, also referred to as the giant component. The first metric, robustness, measures the fraction of nodes remaining in the largest connected component as the grpah is being perturbed during anonymization. The second utility metric considers the task of community detection and therewith looks at the meso-level structure of the graph. Nodes in networks often tend to form communities which can be found by using community detection algorithms [61]. Since randomization is involved in these algorithms, more stable communities can be found by adding a consensus clustering step [62]. To determine how well the community structure is preserved, we compute the normalized mutual information (NMI) [62] between the communities found before and after anonymization. Third, nodes can have different roles in the network and some nodes have a more central position than others. This can be measured by centrality measures such as betweenness centrality, which measures the fraction of shortest paths going through a given node [63]. To determine how well the most central nodes are preserved, we compute the overlap in the top 100 most central nodes before and after anonymization.
@@ 3.2 $k$-Anonymity and equivalence
In order for a node to be $k$-anonymous, it should be equivalent to at least $k-1$ other nodes according to a particular anonymity measure $M$ (discussed in Section 4). Any equivalence measure $M$ partitions the set of nodes into a set of equivalence classes $P_{M}$. If two nodes $v,w\in V$ are equivalent using a measure $M$, we denote this as $v\cong_{M}w$. $P_{M}(v)$ denotes the equivalence class of a given node $v$. A node $v$ is $k$-anonymous when $|P_{M}(v)|\geq k$. We summarize the anonymity of a graph given a certain measure by its uniqueness, a commonly used measure [11] which is determined using the formula in Equation 1. It is essentially the inverse of anonymity and equals the fraction of nodes for which there is no other equivalent node, i.e., $k=1$.
$U_{M}(G)=\frac{|\{v:v\in V,|P_{M}(v)|=1\}|}{|V|}$ (1)
4 $k$-Anonymity measures
In this section, we focus on $k$-anonymity, as previously discussed in Section 2.4, and give an overview of measures for $k$-anonymity introduced in literature. We categorize these based on what type of information they take into account: 1) degree based, 2) neighborhood based, 3) automorphism based and 4) hybrid measures. For each measure we describe both the measure itself and, even though the anonymization process itself is beyond the scope of this work, give an estimate of how complex it is to, given this measure, achieve anonymity with as few alterations as possible. A parameterized version for most measures can be created to account for structural information beyond the direct neighborhood of the considered node, denoted by parameter $d$. Additionally, we discuss how the measures can be extended by a cascading step. For completeness, we end the overview by summarizing measures that do not fit in the four categories. Note that while various works have introduced measures for network extensions such as node labels [64, 65, 66, 67], edge labels [68] or edge weights [69], in the remainder of this work we focus on undirected unlabeled networks.
Table 2 summarizes the measures we focus on in the remainder of this paper. These measures are chosen based on two criteria. First, each of the measures assigns one or more values to each node, which are used to divide the set of all nodes into equivalence classes. Second, the chosen measures model feasible attacker scenarios.
@@ 4.1 Degree based
We distinguish between three measures within the category of degree based measures, being degree, vertex refinement queries and a measure based on joint degree.
@@
Most of the works using degree based measures focus on the degree of the node [31, 78, 79, 80, 81, 70, 71, 72]. A graph is then $k$-anonymous if each degree occurs at least $k$ times in the network. Compared to the other measures we discuss, this is the simplest measure, as it accounts for the least structural information and is trivial to compute for a given node. As a result, anonymization algorithms have been introduced that, given this measure, use the minimum number of edge deletions and additions to obtain an anonymous graph. These approaches can be used on networks with over a million nodes and edges [31, 71, 72]. Several works specifically aim to preserve certain network properties to retain utility [79, 81, 70]. The work in [82] builds upon the notion of $k$-degree anonymity and assumes there might be uncertainty in the knowledge of the attacker. This is accounted for by a so-called binning approach in which nodes with a similar degree, which are in the same bin, are also equivalent. Since this approach accounts for uncertainty, it can merge equivalence classes that are distinct according to the degree measure, hence it is as most as strict as this measure.
@@
In  [76, 32], a parameterized measure is introduced that takes into account the degree of nearby nodes is introduced. This work introduces the notion of vertex refinement queries (VRQ), denoted $\mathcal{H}_{i}$, for which the distance can be set with parameter $i$. When $\mathcal{H}_{1}$ is used, only the degree of the node is taken into account. For $\mathcal{H}_{i}$, with $i>1$, the degree distribution of the nodes at distance at most $i-1$ is taken into account. The measure is applied recursively, i.e., $\mathcal{H}_{i}$ is used to compute $\mathcal{H}_{i+1}$ for each $i\geq 1$. This measure is more difficult to anonymize for and requires an approximation algorithm to minimize the number of edge deletions to arrive at an anonymous graph.
@@
A measure that is more strict than degree, but not as strict as VRQ, has been introduced in [83] and accounts for degrees of all node pairs connected by an edge. In this case, a node can be identified if at least one of its edges has a unique combination of endpoint degrees. While this measure accounts for only slightly more information than node degree itself, anonymization is already much more difficult. To anonymize a network given this measure, the work proposes an integer programming problem and an approximation algorithm to anonymize larger networks. This measure differs from the remaining measures as it does not compute a value for a node based on which the equivalence classes are determined. Hence we do not include the measure in our comparison. However, it does show conceptual similarities to the cascading algorithm we discuss in Section 4.5 and 6.3.
4.2 Neighborhood based
@@
For neighborhood based measures, we distinguish between measures that account for complete structural information and measures that are based on meaningful properties of the neighborhood. The most commonly used neighborhood based measure accounts for the complete structure of the 1-neighborhood [11, 74, 64, 65, 66, 75]. Here nodes are said to be equivalent if their neighborhoods are isomorphic, which implies the neighborhoods are not distinguishable based on network structure. Other works propose a parameterized version of this measure assuming structural knowledge of the $d$-neighborhoods [75, 66, 73, 12], resulting in a measure referred to as $d$-$k$-anonymity. Similarly to VRQ, this measure is computed recursively. Hence, in order for a node to be $d$-$k$-anonymous, with $d\geq 2$, it needs to be $(d-1)$-$k$-anonymous.
@@
The work in [73] uses heuristics to speed up the computation of $d$-$k$-anonymity. These heuristics are graph invariants: properties that are equal if two graphs, or in this case $d$-neighborhoods, are isomorphic. If the values are not equal, the $d$-neighborhoods can not be isomorphic and as a result the nodes can not be equivalent. However, the converse does not hold: if the graph invariants are equal it does not imply that the neighborhoods are isomorphic. The heuristics used are count, which measures the number of nodes and edges in the $d$-neighborhood and degrees, which we will refer to as degdist, which measures the degree distribution of the $d$-neighborhood. As these heuristics in themselves can be used as measures for equivalence and are in terms of what information they measure elegantly positioned between degree and $d$-$k$-anonymity, we choose to include these in our comparison.
4.3 Hybrid